Use the LaPlace Transform to solve the given IVP.

y′′ + 4 y= -10e^−t y(=0) 0,−=y′(0) 4
Answer #1

St. g +4y = -10 *, y 10) 201410)=4 => [ {y+)3 + 4 + {y(t)} = -10%{ety 22sY (5) - SY(0)- ylo) + 4/19 -20 » s’ycs) ++ ++y15)=> -48-14 E ACs+4) +(BS+c) (5+1) 334,4-14 = f() =) ^= -405-2 : :: . A = -2 coefficients of s², o = A TB B=2 (oefficients of s